Accelerator Update for May 21, 2008 - May 23, 2008
The day shift began with Operations monitoring Tevatron (TeV) store 6153 (85.31E30) and the Recycler stash (297.1E10), with the Antiproton Source (Pbar) stacking (29.7E10), with the Main Injector, Booster, and Linac operating normally, with the Collider Detector at Fermilab (CDF) and DZero (D0) using the colliding beams, and with the NuMI, SciBooNE / MiniBooNE (BNB), and the Meson MTest experiments taking beam.
An expert reported at 9:40 AM that the TeV sector F2 BLM system was operating normally.
Operations received a report at 10 AM that a MI power supply (I:V901V) had a LCW leak. Operators verified the leak and then contacted a power supply expert.
A controls expert reported at 10:03 AM that she was investigating the Recycler VSA problems.
MI experts completed their study at noon.
At 12:25 PM, KRF3 tripped off. It reset.
A power supply expert reported at 2:30 PM that he was swapping out a Booster quadrupole power supply (B:QL3).
At 2:53 PM, KRF2 tripped off. It reset.
Operations reported at 4 PM that the Recycler had stashed antiprotons three times during the day shift.
A Recycler expert explained to Operations at 4:22 PM how to determine the beam density if the VSA was not working.
At 5:54 PM, KRF6 tripped off. It reset.
Operations disabled the MTest beam at 6:09 PM.
At 6:50 PM, Operations reported that MI beam was aborting due to bad Booster beam. Operators lowered the Booster beam intensity and then slowly brought it back up to normal.
At 6:55 PM, KRF4 tripped off twice in 30 minutes. It reset.
Per the Run Coordinator, Operations halted stacking at 9 PM in preparation for the scheduled target replacement.
Thursday May 22
The midnight shift began with Operations monitoring store 6153 (40.29E30), the stash (369.1E10), and the stack (160.6E10), with Pbar in standby, and with NuMI and BNB taking beam.
At 1:17 AM, KRF6 tripped off. It reset.
Operations terminated store 6153 at 1:49 AM.
Operations put the TeV into shot setup at 2 AM.
At 2:10 AM, KRF2 tripped off. It reset.
Operations established store 6154 at 3:31 AM with an initial luminosity of 278.2E30.
Recycler experts began an Electron Cooling drag rate measurement around 3:38 AM.
At 3:44 AM, KRF6 tripped off and would not reset. Operators had to adjust the CS switch balance to reset the station.
At 4:47 AM, KRF4 tripped off. It reset.
Operations prepared the Pbar Pretarget area for access at 6:04 AM. Experts will begin replacing the antiproton target at 7 AM.
Recycler experts completed their measurements at 7:31 AM.
Operations reported at 8:39 AM that the NuMI LCW leak was found on a manifold on the upstream end of Q120. A NuMI expert said that it would be an easy repair.
Operations reported at 11:37 AM that operators had search and secured the NuMI decay pipe tunnel and the Absorber Hall.
A MINOS experts reported that the Far Detector work was completed at 12:10 PM.
At 12:16 PM, Linac experts reported that the plate supply for LRF3 had failed. They began further investigations.
At 12:17 PM, Operations reported that they had lost control over the MI-50 pond water pumps. Operations contacted experts.
A MI expert began making stacktail measurements at 12:56 PM.
Linac came out of access at 1:59 PM.
Operators completed their search and secure of the NuMI target hall at 2 PM.
Operations established beam in Linac at 3 PM.
Booster and MI came out of access at 3:11 PM.
A Pbar expert reported that the new antiproton target was ready for beam at 3:17 PM.
Operations established beam to Pbar, NuMI, and BNB at 3:53 PM.
At 4:06 PM, LRF2 tripped off for the first of two times in 8 minutes. It reset.
At 5:42 PM, KRF6 tripped off. It reset.
At 5:49 PM, the NuMI beam permit tripped off due to the current drift of a power supply (E:V108). Everything reset.
A Pbar expert began conducting a target scan at 5:56 PM.
At 6:02 PM, the NuMI beam permit tripped off due to an incomplete RF gap envelope. MIRF16 was the problem station; operators bypassed the station.
Operations established beam to MTest at 6:56 PM.
At 7:02 PM, KRF3 tripped off. It reset.
At 7:37 PM, LRF4 tripped off. It reset.
A MI expert began a coalescing study at 8:27 PM.
Operations reported at 11:58 PM that the Recycler had stashed four times during the evening shift.
Operations disabled the MTest beam at midnight.
The midnight shift began with Operations monitoring store 6154 (54.34E30) and the stash (232.2E10), with Pbar stacking (38.4E10), and with NuMI and BNB taking beam.
At 12:25 AM, KRF6 tripped off. It reset.
At 1:08 AM, store 6154 aborted and the TeV ramp dumped. Operators reported that there was a shunt trip at C0. Operators investigated and discovered that the problem was with a Standby Power Unit (SPU) at TeV sector D3. Operations called in an expert.
At 1:44 AM, a MI correction magnet power supply (I:H502) tripped off. Operators found and replaced a bad power supply. They also replaced its regulator card, but the supply would not come on. Operations contacted an expert who said that a breaker had probably tripped off and he would come in to reset it.
At 1:58 AM, the Booster East Anode Power Supply tripped off and took down BRF2. Everything reset.
Operations put Pbar into standby at 2:21 AM.
At 2:43 AM, BRF7 & 8 tripped off on a bad vacuum indication even though no vacuum pumps had tripped off. Everything reset.
At 3:08 AM, KRF3 tripped off. It reset.
A power supply expert reported at 3:34 AM that he had brought I:H502 online. Pbar resumed stacking.
At 3:48 AM, BRF19 tripped off. It reset.
At 4:16 AM, Operations reported that problems with beam at the MI-8 collimators required them to lower the beam intensity for BNB.
At 4:46 AM, KRF6 tripped off. It reset.
At 4:48 AM, KRF5 tripped off. It reset.
At 5:11 AM the TeV ramp dumped due to a D2 power supply going “Not Ready.”
At 5:24 AM, KRF6 tripped off. It reset.
Operations established beam to MTest at 6:05 AM.
A power supply expert reported at 6:16 AM that he had replaced the SPU at TeV sector D2.
Operations put the TeV through a dry squeeze at 7:22 AM.
At 7:28 AM, KRF6 tripped off. It reset.
Operations reported at 8 AM that the Recycler had stashed four times during the midnight shift.
The Plans for Friday and the Weekend
The plans for today are to repair a problem with the TeV sector C4 cold compressor, to allow the water experts to makeup water for the TeV LCW system, to let Booster experts tune, and then to resume normal stacking, stashing, and storing for the weekend.
Comments and Suggestions
|last modified 05/23/2008 email Fermilab|
|Security, Privacy, Legal|